# Performance bounds for priority-based stochastic coflow scheduling

### v1

@@ 3.4 Performance guarantees for the Sincronia-based priority policy
Let $\pi$ be the priority order computed by Algorithm 1 from expected processing times. We now consider the greedy rate allocation of $\mathtt{Sincronia}$(see Algorithm 2 in [3]). We let $C_{k}$ be the random completion time of coflow $k$ under the greedy rate allocation when coflows are processed in the priority order $\pi$. It is proven in [3] that $c_{\pi(t)}\leq 2\max_{\ell}\sum_{\tau=1}^{t}p_{\ell,\pi(\tau)}$ for any fixed realization of flow volumes, so that
@@ 4.1 Distribution-independent upper bound
For a given priority order $\theta$, let the matrix $\mathbf{P}^{\theta}_{k}=[P_{\ell,j}]$ denote the $L_{k}\times k$ matrix with the processing times of flows ahead of coflow $k$ on the $L_{k}$ links used by coflow $k$ in the order $\theta$. For ease of notation, we use $k$ instead of $\theta(k)$. It will be apparent from the context that $\mathbf{P}^{\theta}_{k}$ denotes the $\mathbf{P}$ matrix for the $k$th coflow in the order $\theta$ which is in fact $\theta(k)$. Note that the matrix $\mathbf{P}^{\theta}_{k}$ can be written as $\mathbf{P}^{\theta}_{k}=\mathbf{M}^{\theta}_{k}+\mathbf{X}^{\theta}_{k}$, where $\mathbf{M}^{\theta}_{k}=\mathbb{E}\mathbf{P}^{\theta}_{k}$ and $\mathbf{X}^{\theta}_{k}$ is a zero-mean matrix. We observe from (11) that
@@
[…]
@@
Thus, computing an upper-bound on $\alpha$ can be reduced to the problem of computing the expectation of the $\infty$-norm of random matrices whose elements are taken from the given distribution. These bounds will depend on the probability distribution of $X_{\ell,j}$s. For a finite $k$ but $L$ going to $\infty$, one can use extreme-value theory [22] which models how the maximum of a certain number of random variables behaves. When both $k$ and $L$ go to $\infty$, we refer the reader to [1] for results on the asymptotic bounds on $\alpha$. Below, we give an upper bound on $\alpha$ that is non-asymptotic but less sharp and which only depends upon the existence of certain moments.
Let $\mu_{min}=\min\{\mu_{\ell,j}:\mu_{\ell,j}>0\}$ and let $p\in(1,\infty)$ be such that $m_{p}:=\max_{\ell,j}E|X_{\ell,j}|^{p}<\infty$. That is, $\mu_{min}$ is the smallest expected flow-size among those flows that are not always 00, and $m_{p}$ is the largest $p$th moment of the flow-sizes provided that it is finite.
@@
Let $\mu_{min}=\min_{\ell,j}\mu_{\ell,j}$ and let $m_{p}:=\max_{\ell,j}E|X_{\ell,j}|^{p}<\infty$ for $p\in(1,\infty)$. Then,
@@
Let $Y_{i}$, $i=1,2,\ldots,n$ be independent random variables with $E|Y_{i}|^{p}\leq m_{p},\forall i$. Then
@@
The result is a consequence of the Cauchy-Schwartz inequality which applied on $\sum_{i}|Y_{i}|$ gives
@@
Now taking expectation on both sides and using the linearity of the expectation operator and the fact that $\mathbb{E}|Y_{i}|^{p}\leq m_{p}$, we obtain the claimed result: ∎
@@
$\leq\frac{\left(\sum_{\ell}k^{p}m_{p}\right)^{p}}{k\mu_{min}}=L^{1/p}\frac{m_{p}^{1/p}}{\mu_{min}},$ (21)
where (20) follows from Jensen’s inequality and linearity of expectation, and (21) follows from Lemma 1 and the fact that $E|X_{\ell,j}|^{p}=m_{p}$.
@@
$\alpha\leq 1+L^{1/2}\frac{\sigma_{max}}{\mu_{min}}$
@@
This bound in Theorem 4 is consistent with $\mathtt{Sincronia}$in the sense that if the flow sizes are deterministic, then $|X|^{p}=0$ for all $p>1$, and we recover the approximation factor of $\mathtt{Sincronia}$in the clairvoyant setting.
4.2 Asymptotically matching asymptotic lower bound
To show that the upper bound in Theorem 4 is asymptotically tight, we shall assume that $P_{\ell,j}$ are Pareto random variables and give asymptotic bounds as $L_{k}\to\infty$.
@@ 4.2 Asymptotically matching asymptotic lower bound
A Pareto random variable of shape parameter $\zeta$ belongs to the class of regularly varying random variables of index $\zeta$ (see [20] and references therein for the definition and properties of the class of regularly varying random variables).
From the closure property of regularly varying random variables [20], $\sum_{j=1}^{k}P_{\ell,j}$ is again a regularly varying random variable of index $\zeta$ (which is the shape parameter of the constituent Pareto random variables). Further, both $P_{\ell,j}$ and $\sum_{j=1}^{k}P_{\ell,j}$ belong to the maximum domain of attraction of the Fréchet distribution of index $\zeta$.
For this class of distributions, the maximum grows as $L^{1/\zeta}$ [20], that is,
@@ 4.3 Distribution-dependent tighter upper bounds
See appendix B
Combining Theorem 3 and Lemma 2, we have the following result regarding the approximation ratio of $\mathtt{Sincronia}$for gamma-distributed processing times.
@@
If all random variables $P_{\ell,j}$ are independent and gamma-distributed, the approximation of $\mathtt{Sincronia}$with expected processing times as inputs is $4(c+\log\left(c+\sqrt{c^{2}-1}\right)0$.
@@
This bound on $\alpha$ does not directly depend upon $CV_{max}$. However, if $CV_{max}=0$, then $\gamma=1$ which leads to $\alpha=1$. We thus recover the approximation ratio of $4$ of $\mathtt{Sincronia}$in the clairvoyant setting.
@@
If all random variables $P_{\ell,j}$ are independent and follow a normal distribution with mean $\mu_{\ell,j}$ and coefficient of variation $CV_{\ell,j}$, then $\alpha\leq 1+\sqrt{2\log(L)}CV_{max}$, where $L=\left|{\mathcal{L}}\right|$ is the number of ports and $CV_{max}=\max_{\ell,j}CV_{\ell,j}$.
@@
If all random variables $P_{\ell,j}$ are independent and and follow a normal distribution, the approximation ratio of the $\mathtt{Sincronia}$-based priority policy is $4\left\{1+\sqrt{2\log(L)}CV_{max}\right\}$.
Note in particular that when $CV_{max}=0$, we retrieve the approximation ratio of $\mathtt{Sincronia}$in the clairvoyant setting. Further, the upper bound only grows as $\Theta(\sqrt{\log(L)})$ which is even slower than the $\Theta(\log(L)$ growth for the gamma distribution. This is to be expected since the normal distribution has a lighter tail than the gamma distribution.
@@ 5 Numerical Results
The random instances are generated using a coflow workload generator based on a real-world trace with $526$ coflows. The trace was obtained from a one-hour long run of Hive/MapReduce jobs on a $3000$-machine cluster with $150$ racks at Facebook [12]. The workload generator allows upsampling the Facebook trace to desired number of ports $L$ and coflows $n$, while still keeping workload characteristics similar to the original Facebook trace. The original workload generator, which is publicly available [2], was slightly modified in order to be able to generate flow sizes following a target probability distribution (e.g., a normal or Gamma distribution with specified mean and standard deviation).
@@
  • The number of ports is always assumed to be even (the first $L/2$ ports are input ports, while the other ports are output ports) and can take values between $8$ and $80$.
  • The number of coflows $n$ is either $n=L/2$, $n=L$ or $n=3\,L/2$.
  • When the flow sizes follow a specific target distribution, the mean flow size is $100$ and the coefficient of variation $CV$ is either $\frac{1}{2}$, $1$ or $2$. Otherwise, flow sizes are drawn from the empirical distribution of the Facebook trace.
@@ 5.1 Small instances
Let us first consider the results obtained with a Gamma distribution for flow volumes. We consider three different coefficients of variations: $CV=\frac{1}{2}$, $CV=1$ and $CV=2$.
The mean, standard deviation as well as the first and third quartiles of the ratios $C^{ALG}/C^{LP}$ are reported in Tables 1, 2 and 3 for flow sizes following a Gamma distribution with $CV=0.5,1$, and $2$, respectively. For $CV=0.5$, we observe that for all considered configurations of the numbers of ports and coflows, the non-clairvoyant Sincronia performs almost as well as the clairvoyant one. We also note that the variability of the ratio $C^{ALG}/C^{LP}$ is also quite low for both algorithms, indicating that they perform very well for the vast majority of our random problem instances. We also observe that the ratio $C^{RR}/C^{LP}$ has a much higher mean and standard deviation, which shows that the round-robin allocation performs poorly as compared to our NC priority policy when the variability of flow sizes is low. When the variability of the flow sizes increases (i.e., $CV=1$ or $CV=2$), the performance of the non-clairvoyant Sincronia degrades as compared to the clairvoyant one. The variability of the ratio $C^{NC}/C^{LP}$ is also much higher that that observed in the case $CV=\frac{1}{2}$. Nevertheless, the performance of the NC remains fairly close to that of CL, and much better than that obtained with RR.
CL NC RR
@@
4 4 1.01 0.06 0.97 1.06 1.08 0.06 1.02 1.11 1.55 0.12 1.48 1.67
8 4 1.07 0.06 1.03 1.11 1.11 0.08 1.06 1.14 1.71 0.19 1.60 1.83
8 8 1.05 0.05 1.01 1.08 1.09 0.05 1.06 1.13 2.08 0.17 2.01 2.19
16 8 1.09 0.04 1.06 1.12 1.11 0.05 1.08 1.15 2.23 0.18 2.10 2.34
16 16 1.09 0.03 1.06 1.11 1.12 0.05 1.08 1.15 2.80 0.20 2.66 2.94
32 16 1.10 0.03 1.08 1.13 1.12 0.04 1.08 1.15 2.92 0.24 2.78 3.03
Table 1: $C^{ALG}/C^{LP}$ for flow sizes following a Gamma distribution with $CV=0.5$.
CL NC RR
@@
4 4 0.96 0.09 0.90 1.02 1.14 0.09 1.09 1.20 1.38 0.15 1.28 1.49
8 4 1.10 0.10 1.03 1.18 1.20 0.11 1.13 1.28 1.63 0.21 1.50 1.78
8 8 1.04 0.08 1.00 1.10 1.18 0.08 1.14 1.23 1.90 0.19 1.77 2.03
16 8 1.13 0.06 1.10 1.18 1.21 0.07 1.17 1.26 2.17 0.19 2.05 2.29
16 16 1.12 0.04 1.09 1.15 1.20 0.06 1.16 1.24 2.65 0.18 2.51 2.76
32 16 1.14 0.04 1.11 1.18 1.18 0.05 1.15 1.22 2.84 0.22 2.70 2.95
Table 2: $C^{ALG}/C^{LP}$ for flow sizes following a Gamma distribution with $CV=1$.
CL NC RR
@@
4 4 0.89 0.11 0.81 0.98 1.28 0.13 1.21 1.36 1.14 0.17 1.02 1.27
8 4 1.18 0.18 1.03 1.31 1.43 0.17 1.33 1.52 1.54 0.27 1.35 1.74
8 8 1.05 0.12 0.96 1.13 1.41 0.12 1.32 1.50 1.63 0.23 1.47 1.80
16 8 1.26 0.10 1.19 1.33 1.49 0.10 1.41 1.55 2.09 0.23 1.93 2.25
16 16 1.20 0.07 1.16 1.24 1.46 0.08 1.40 1.51 2.40 0.19 2.27 2.55
32 16 1.27 0.07 1.21 1.32 1.40 0.08 1.35 1.45 2.76 0.22 2.60 2.89
Table 3: $C^{ALG}/C^{LP}$ for flow sizes following a Gamma distribution with $CV=2$.
For a more detailed analysis, we also plot the histograms of the ratios $C^{ALG}/C^{LP}$ obtained for the $100$ random problem instances in the case $L=32$ and $n=16$ in Figure 1(a) and Figure 1(b) for $CV=0.5$ and $CV=2$, respectively. For $CV=0.5$, clearly the histograms obtained for $ALG=CL$ and $ALG=NC$ are very similar. The values obtained with RR are much higher, and even higher than those obtained with a random priority order. For $CV=2$, as compared to the histogram for $CL$, the most probable values for $NC$ are slightly shifted to the right by $+0.15$. As in the case $CV=\frac{1}{2}$, the values obtained with RR are much higher, and even higher than those obtained with a random priority order.
(a) $CV=0.5$
(b) $CV=2$
Figure 1: Histograms of the ratios $C^{ALG}/C^{LP}$ for CL, NC, RR and a random-order policy (RO) for $100$ random problem instances with $L=32$ ports and $n=16$ coflows. Flow sizes follow a Gamma distribution with either (a) $CV=0.5$ or (b) $CV=2$.
Figure 2 summarizes the above results. It shows the average values (over $100$ random instances) obtained for CL, NC, and RR. We first note that CL provides near-optimal results in all cases. For $CV=0.5$, NC performs almost as well as CL and much better than RR. For $CV=2$, there is a noticeable performance degradation as compared to CL, but the proposed algorithm still outperforms significantly RR. We also note that the gap between CL and NC decreases as the number of ports increases.
Figure 2: Average values of $C^{ALG}/C^{LP}$ for flow sizes following a Gamma distribution with $CV=0.5$ (left), $CV=1$ (middle) or $CV=2$ (right). The number of ports $L$ and the number $n$ of coflows are indicated on the $x$-axis as $L.n$.
@@
We now consider the results obtained for normally-distributed flow volumes. As before, we consider three different coefficients of variations: $CV=\frac{1}{2}$, $CV=1$ and $CV=2$. The mean, standard deviation as well as the first and third quartiles of the ratios $C^{ALG}/C^{LP}$ are reported in Tables 4, 5 and 6 for $CV=\frac{1}{2}$, $CV=1$ and $CV=2$, respectively.
As for the gamma distribution, we observe that when the variability of flow sizes is low the NC performs almost as well as CL and provides almost optimal solutions. We also observe that the variability of the ratio $C^{NC}/C^{LP}$ is quite low in all configurations and that NC provides much better results than RR. These observations are confirmed by the histograms shown in Figure 3(a) which report the ratios $C^{ALG}/C^{LP}$ obtained for the $100$ random problem instances in the case $L=32$ and $n=16$. Although the repartition is slightly different, CL and NC provide similar results, which are much better than those obtained with RR or RO.
CL NC RR
@@
4 4 1.02 0.06 0.98 1.08 1.09 0.06 1.04 1.12 1.58 0.13 1.51 1.69
8 4 1.08 0.06 1.04 1.13 1.12 0.08 1.06 1.15 1.74 0.19 1.62 1.86
8 8 1.05 0.05 1.02 1.09 1.10 0.05 1.06 1.13 2.12 0.18 2.03 2.23
16 8 1.09 0.04 1.07 1.12 1.12 0.05 1.08 1.15 2.27 0.18 2.14 2.36
16 16 1.09 0.03 1.07 1.11 1.12 0.05 1.08 1.15 2.84 0.20 2.70 2.98
32 16 1.11 0.04 1.08 1.13 1.12 0.04 1.08 1.15 2.95 0.24 2.80 3.06
Table 4: $C^{ALG}/C^{LP}$ for flow sizes following a Normal distribution with $CV=0.5$.
@@
4 4 1.08 0.09 1.01 1.13 1.22 0.08 1.17 1.27 1.58 0.17 1.50 1.69
8 4 1.20 0.09 1.13 1.25 1.27 0.10 1.19 1.33 1.83 0.22 1.69 1.98
8 8 1.14 0.07 1.10 1.19 1.24 0.07 1.20 1.29 2.16 0.21 2.05 2.30
16 8 1.21 0.05 1.17 1.25 1.26 0.07 1.22 1.30 2.40 0.20 2.27 2.53
16 16 1.20 0.04 1.17 1.22 1.26 0.06 1.22 1.29 2.98 0.21 2.83 3.13
32 16 1.22 0.04 1.19 1.25 1.25 0.05 1.21 1.28 3.15 0.25 3.00 3.27
Table 5: $C^{ALG}/C^{LP}$ for flow sizes following a Normal distribution with $CV=1$.
CL NC RR

### v2

@@ 3.4 Performance guarantees for the Sincronia-based priority policy
Let $\pi$ be the priority order computed by Algorithm 1 from expected processing times. We now consider the greedy rate allocation of $\mathtt{Sincronia}\$(see Algorithm 2 in [Agarwal2018]). We let $C_{k}$ be the random completion time of coflow $k$ under the greedy rate allocation when coflows are processed in the priority order $\pi$. It is proven in [Agarwal2018] that $c_{\pi(t)}\leq 2\max_{\ell}\sum_{\tau=1}^{t}p_{\ell,\pi(\tau)}$ for any fixed realization of flow volumes, so that
@@ 4.1 Distribution-independent upper bound
For a given priority order $\theta$, let the matrix $\mathbf{P}^{\theta}_{k}=[P_{\ell,j}]$ denote the $L_{k}\times k$ matrix with the processing times of flows ahead of coflow $k$ on the $L_{k}$ ports used by coflow $k$ in the order $\theta$. For ease of notation, we use $k$ instead of $\theta(k)$. It will be apparent from the context that $\mathbf{P}^{\theta}_{k}$ denotes the $\mathbf{P}$ matrix for the $k$th coflow in the order $\theta$ which is in fact $\theta(k)$. Note that the matrix $\mathbf{P}^{\theta}_{k}$ can be written as $\mathbf{P}^{\theta}_{k}=\mathbf{M}^{\theta}_{k}+\mathbf{X}^{\theta}_{k}$, where $\mathbf{M}^{\theta}_{k}=\mathbb{E}\mathbf{P}^{\theta}_{k}$ and $\mathbf{X}^{\theta}_{k}$ is a zero-mean matrix. We observe from (11) that
@@
where $\|\cdot\|$ is the induced $\infty$-norm and $\pi$ is the order computed by Sincronia. An upper bound for $\alpha$ can be obtained by taking the worst-case order and using the triangular inequality for norms,
@@
Thus, computing an upper-bound on $\alpha$ can be reduced to the problem of computing the expectation of the $\infty$-norm of random matrices whose elements are taken from the given distribution. These bounds will depend on the probability distribution of $X_{\ell,j}$s. For a finite $k$ but $L$ going to $\infty$, one can use extreme-value theory [EKM11] which models how the maximum of a certain number of random variables behaves. When both $k$ and $L$ go to $\infty$, we refer the reader to [APS23] for results on the asymptotic bounds on $\alpha$. Below, we give an upper bound on $\alpha$ that is non-asymptotic but less sharp and which only depends upon the existence of certain moments.
Let $\mu_{min}=\min\{\mu_{\ell,j}:\mu_{\ell,j}>0\}$ and let $p\in(1,\infty)$ be such that $m_{p}:=\max_{\ell,j}\mathbb{E}|X_{\ell,j}|^{p}<\infty$. That is, $\mu_{min}$ is the smallest expected flow-size among those flows that are not always 0, and $m_{p}$ is the largest $p$th moment of the flow-sizes provided that it is finite.
@@
Let $\mu_{min}=\min_{\ell,j}\mu_{\ell,j}$ and let $m_{p}:=\max_{\ell,j}\mathbb{E}|X_{\ell,j}|^{p}<\infty$ for $p\in(1,\infty)$. Then,
@@
Let $Y_{i}$, $i=1,2,\ldots,n$ be independent random variables with $\mathbb{E}|Y_{i}|^{p}\leq m_{p},\forall i$. Then
@@
The result is a consequence of the Hölder inequality which applied on $\sum_{i}|Y_{i}|$ gives
@@
Now taking expectation on both sides and using the linearity of the expectation operator and the fact that $\mathbb{E}|Y_{i}|^{p}\leq m_{p}$, we obtain the claimed result. ∎
@@
$\leq\frac{\left(\sum_{\ell}k^{p}m_{p}\right)^{1/p}}{k\mu_{min}}=L^{1/p}\frac{m_{p}^{1/p}}{\mu_{min}},$ (21)
where (20) follows from Jensen’s inequality and linearity of expectation, and (21) follows from Lemma 1 and the fact that $\mathbb{E}|X_{\ell,j}|^{p}=m_{p}$.
@@
$\alpha\leq 1+L^{1/2}\frac{\sigma_{max}}{\mu_{min}}$ (22)
@@
This bound in Theorem 4 is consistent with $\mathtt{Sincronia}\$in the sense that if the flow sizes are deterministic, then $|X|^{p}=0$ for all $p>1$, and we recover the approximation factor of $\mathtt{Sincronia}\$in the clairvoyant setting.
4.2 Asymptotically matching lower bound
To show that the upper bound in Theorem 4 is asymptotically tight, we shall assume that flow volumes are Pareto random variables and give asymptotic bounds as $L_{k}\to\infty$.
In this section as well as in Section 4.3, to simplify the notation, we shall assume, without loss of generality that the order $\theta$ is the ascending order. The results hold for any fixed order $\theta$ since the distribution of the flow volumes are independent of the order and flow volumes will be assumed to be independent and identically distributed.
@@ 4.2 Asymptotically matching lower bound
A Pareto random variable of shape parameter $\zeta$ belongs to the class of regularly varying random variables of index $\zeta$ (see [Mik99] and references therein for the definition and properties of the class of regularly varying random variables).
From the closure property of regularly varying random variables [Mik99], $\sum_{j=1}^{k}P_{\ell,j}$ is again a regularly varying random variable of index $\zeta$ (which is the shape parameter of the constituent Pareto random variables). Further, both $P_{\ell,j}$ and $\sum_{j=1}^{k}P_{\ell,j}$ belong to the maximum domain of attraction of the Fréchet distribution of index $\zeta$.
For this class of distributions, the maximum grows as $L^{1/\zeta}$ [Mik99], that is,
@@ 4.3 Distribution-dependent tighter upper bounds
See Appendix B. ∎
Combining Theorem 3 and Lemma 2, we have the following result regarding the approximation ratio of $\mathtt{Sincronia}\$for gamma-distributed processing times.
@@
If all random variables $P_{\ell,j}$ are independent and gamma-distributed, the approximation of $\mathtt{Sincronia}\$with expected processing times as inputs is $4(c+\log\left(c+\sqrt{c^{2}-1}\right))$.
@@
This bound on $\alpha$ does not directly depend upon $\eta_{max}$, which is an upper bound on the coefficient of variation of the processing times. However, if $\eta_{max}=0$, then $\gamma=0$ which leads to $\alpha=1$. We thus recover the approximation ratio of $4$ of $\mathtt{Sincronia}\$in the clairvoyant setting.
@@
If all random variables $P_{\ell,j}$ are independent and follow a normal distribution with mean $\mu_{\ell,j}$ and coefficient of variation $\eta_{\ell,j}$, then $\alpha\leq 1+\sqrt{2\log(L)}\eta_{max}$, where $L=\left|{\mathcal{L}}\right|$ is the number of ports and $\eta_{max}=\max_{\ell,j}\eta_{\ell,j}$.
@@
If all random variables $P_{\ell,j}$ are independent and and follow a normal distribution, the approximation ratio of the $\mathtt{Sincronia}$-based priority policy is $4\left\{1+\sqrt{2\log(L)}\eta_{max}\right\}$.
Note in particular that when $\eta_{max}=0$, we retrieve the approximation ratio of $\mathtt{Sincronia}\$in the clairvoyant setting. Further, the upper bound only grows as $\Theta(\sqrt{\log(L)})$ which is even slower than the $\Theta(\log(L)$ growth for the gamma distribution. This is to be expected since the normal distribution has a lighter tail than the gamma distribution.
@@ 5 Numerical Results
The random instances are generated using a coflow workload generator based on a real-world trace with $526$ coflows. The trace was obtained from a one-hour long run of Hive/MapReduce jobs on a $3000$-machine cluster with $150$ racks at Facebook [Chowdhury2014]. The workload generator allows upsampling the Facebook trace to desired number of ports $L$ and coflows $n$, while still keeping workload characteristics similar to the original Facebook trace. The original workload generator, which is publicly available [Sin-Rep], was slightly modified in order to be able to generate flow sizes following a target probability distribution (e.g., a normal or Gamma distribution with specified mean and standard deviation).
@@
  • The number of ports is always assumed to be even (the first $L/2$ ports are input ports, while the other ports are output ports) and can take values between $4$ and $40$.
  • When the flow sizes follow a specific target distribution, the mean flow size is $10$. The coefficient of variation will take different values depending on the size of the instance and will be specified in the relevant place. Otherwise, flow sizes are drawn from the empirical distribution of the Facebook trace.
@@ 5.1 Small instances
Let us first consider the results obtained with a Gamma distribution for flow volumes. We consider three different coefficients of variations: $\eta=0.5$, $\eta=1$ and $\eta=2$.
The mean, standard deviation as well as the first and third quartiles of the ratios $C^{ALG}/C^{LP}$ are reported in Tables 1, 2 and 3 for flow sizes following a Gamma distribution with $\eta=0.5,1$, and $2$, respectively. Column UB in these tables shows the theoretical upper bound on $C^{NC}/C^{LP}$ computed using (23). For $\eta=0.5$, we observe that for all considered configurations of the numbers of ports and coflows, non-clairvoyant Sincronia performs almost as well as the clairvoyant one. We also note that the variability of the ratio $C^{ALG}/C^{LP}$ is also quite low for both algorithms, indicating that they perform very well for the vast majority of our random problem instances. We also observe that the ratio $C^{RR}/C^{LP}$ has a much higher mean and standard deviation, which shows that the round-robin allocation performs poorly as compared to our NC priority policy when the variability of flow sizes is low. When the variability of the flow sizes increases (i.e., $\eta=1$ or $\eta=2$), the performance of the non-clairvoyant Sincronia degrades as compared to the clairvoyant one. The variability of the ratio $C^{NC}/C^{LP}$ is also much higher that that observed in the case $\eta=0.5$. Nevertheless, the performance of NC remains fairly close to that of CL, and much better than that obtained with RR. For all three values of $\eta$, the upper bound UB is pessimistic with values much bigger than $4$ while the empirical ratios are between $1$ and $2$.
UB CL NC $R$
@@
4 4 8.6 1.01 0.06 0.97 1.06 1.08 0.06 1.02 1.11 1.55 0.12 1.48 1.67
8 4 10 1.07 0.06 1.03 1.11 1.11 0.08 1.06 1.14 1.71 0.19 1.60 1.83
8 8 10 1.05 0.05 1.01 1.08 1.09 0.05 1.06 1.13 2.08 0.17 2.01 2.19
16 8 11.2 1.09 0.04 1.06 1.12 1.11 0.05 1.08 1.15 2.23 0.18 2.10 2.34
16 16 11.2 1.09 0.03 1.06 1.11 1.12 0.05 1.08 1.15 2.80 0.20 2.66 2.94
32 16 12.4 1.10 0.03 1.08 1.13 1.12 0.04 1.08 1.15 2.92 0.24 2.78 3.03
Table 1: $C^{ALG}/C^{LP}$ for flow sizes following a Gamma distribution with $\eta=0.5$.
UB CL NC $R$
@@
4 4 15.6 0.96 0.09 0.90 1.02 1.14 0.09 1.09 1.20 1.38 0.15 1.28 1.49
8 4 19.5 1.10 0.10 1.03 1.18 1.20 0.11 1.13 1.28 1.63 0.21 1.50 1.78
8 8 19.5 1.04 0.08 1.00 1.10 1.18 0.08 1.14 1.23 1.90 0.19 1.77 2.03
16 8 23.1 1.13 0.06 1.10 1.18 1.21 0.07 1.17 1.26 2.17 0.19 2.05 2.29
16 16 23.1 1.12 0.04 1.09 1.15 1.20 0.06 1.16 1.24 2.65 0.18 2.51 2.76
32 16 26.6 1.14 0.04 1.11 1.18 1.18 0.05 1.15 1.22 2.84 0.22 2.70 2.95
Table 2: $C^{ALG}/C^{LP}$ for flow sizes following a Gamma distribution with $\eta=1$.
UB CL NC $R$
@@
4 4 36.4 0.89 0.11 0.81 0.98 1.28 0.13 1.21 1.36 1.14 0.17 1.02 1.27
8 4 49.0 1.18 0.18 1.03 1.31 1.43 0.17 1.33 1.52 1.54 0.27 1.35 1.74
8 8 49.0 1.05 0.12 0.96 1.13 1.41 0.12 1.32 1.50 1.63 0.23 1.47 1.80
16 8 61.1 1.26 0.10 1.19 1.33 1.49 0.10 1.41 1.55 2.09 0.23 1.93 2.25
16 16 61.1 1.20 0.07 1.16 1.24 1.46 0.08 1.40 1.51 2.40 0.19 2.27 2.55
32 16 73.0 1.27 0.07 1.21 1.32 1.40 0.08 1.35 1.45 2.76 0.22 2.60 2.89
Table 3: $C^{ALG}/C^{LP}$ for flow sizes following a Gamma distribution with $\eta=2$.
For a more detailed analysis, we also plot the histograms of the ratios $C^{ALG}/C^{LP}$ obtained for the $100$ random problem instances in the case $L=32$ and $n=16$ in Figure 1(a) and Figure 1(b) for $\eta=0.5$ and $\eta=2$, respectively. For $\eta=0.5$, clearly the histograms obtained for $ALG=CL$ and $ALG=NC$ are very similar. The values obtained with RR are much higher, and even higher than those obtained with a random priority order. For $\eta=2$, as compared to the histogram for $CL$, the most probable values for $NC$ are slightly shifted to the right by +0.15+0.15. As in the case $\eta=0.5$, the values obtained with RR are much higher, and even higher than those obtained with a random priority order.
(a) $\eta=0.5$
(b) $\eta=2$
Figure 1: Histograms of the ratios $C^{ALG}/C^{LP}$ for CL, NC, RR and a random-order policy (RO) for $100$ random problem instances with $L=32$ ports and $n=16$ coflows. Flow sizes follow a Gamma distribution with either (a) $\eta=0.5$ or (b) $\eta=2$.
Figure 2 summarizes the above results. It shows the average values (over $100$ random instances) obtained for CL, NC, and RR. We first note that CL provides near-optimal results in all cases. For $\eta=0.5$, NC performs almost as well as CL and much better than RR. For $\eta=2$, there is a noticeable performance degradation as compared to CL, but the proposed algorithm still outperforms significantly RR. We also note that the gap between CL and NC decreases as the number of ports increases.
Figure 2: Average values of $C^{ALG}/C^{LP}$ for flow sizes following a Gamma distribution with $\eta=0.5$ (left), $\eta=1$ (middle) or $\eta=2$ (right). The number of ports $L$ and the number $n$ of coflows are indicated on the $x$-axis as $L.n$.
@@
We now consider the results obtained for normally-distributed flow volumes. As before, we consider three different coefficients of variations: $\eta=\frac{1}{2}$, $\eta=1$ and $\eta=2$. The mean, standard deviation as well as the first and third quartiles of the ratios $C^{ALG}/C^{LP}$ are reported in Tables 4, 5 and 6 for $\eta=\frac{1}{2}$, $\eta=1$ and $\eta=2$, respectively. Column UB in these tables is the theoretical upper bound on $C^{NC}/C^{LP}$ computed using Cor. 3.
As for the gamma distribution, UB is quite pessimistic compared to the observed ratios. Regarding $C^{ALG}/C^{LP}$, we observe that when the variability of flow sizes is low the NC performs almost as well as CL and provides almost optimal solutions. We also observe that the variability of the ratio $C^{NC}/C^{LP}$ is quite low in all configurations and that NC provides much better results than RR. These observations are confirmed by the histograms shown in Figure 3(a) which report the ratios $C^{ALG}/C^{LP}$ obtained for the $100$ random problem instances in the case $L=32$ and $n=16$. Although the repartition is slightly different, CL and NC provide similar results, which are much better than those obtained with RR or RO.
UB CL NC $R$
@@
4 4 7.3 1.02 0.06 0.98 1.08 1.09 0.06 1.04 1.12 1.58 0.13 1.51 1.69
8 4 8.1 1.08 0.06 1.04 1.13 1.12 0.08 1.06 1.15 1.74 0.19 1.62 1.86
8 8 8.1 1.05 0.05 1.02 1.09 1.10 0.05 1.06 1.13 2.12 0.18 2.03 2.23
16 8 8.7 1.09 0.04 1.07 1.12 1.12 0.05 1.08 1.15 2.27 0.18 2.14 2.36
16 16 8.7 1.09 0.03 1.07 1.11 1.12 0.05 1.08 1.15 2.84 0.20 2.70 2.98
32 16 9.3 1.11 0.04 1.08 1.13 1.12 0.04 1.08 1.15 2.95 0.24 2.80 3.06
Table 4: $C^{ALG}/C^{LP}$ for flow sizes following a Normal distribution with $\eta=0.5$.
@@
4 4 10.7 1.08 0.09 1.01 1.13 1.22 0.08 1.17 1.27 1.58 0.17 1.50 1.69
8 4 12.2 1.20 0.09 1.13 1.25 1.27 0.10 1.19 1.33 1.83 0.22 1.69 1.98
8 8 12.2 1.14 0.07 1.10 1.19 1.24 0.07 1.20 1.29 2.16 0.21 2.05 2.30
16 8 13.4 1.21 0.05 1.17 1.25 1.26 0.07 1.22 1.30 2.40 0.20 2.27 2.53
16 16 13.4 1.20 0.04 1.17 1.22 1.26 0.06 1.22 1.29 2.98 0.21 2.83 3.13
32 16 14.5 1.22 0.04 1.19 1.25 1.25 0.05 1.21 1.28 3.15 0.25 3.00 3.27
Table 5: $C^{ALG}/C^{LP}$ for flow sizes following a Normal distribution with $\eta=1$.
UB CL NC $R$